OpenStudy (anonymous):

You want to buy a used car for $8,648.00 plus 4.7% sales tax. You want to make a 5% down payment. How much is your down payment? 10. You want to buy a house in 9 years. The estimated cost is $175,800.00. You want to make a 20% down payment and closing costs are 7%. How much in total do you need save each month to reach your goal if you cover closing costs and the down payment? the sales tax part is messing me up, any step by step help?

OpenStudy (anonymous):

ok so the total cost is House + Tax Total cost = 8648+8648*.047=8648(1+.047)=8648(1.047)=?

OpenStudy (anonymous):

Do you follow?

OpenStudy (anonymous):

not really... why did you add the the car price twice?

OpenStudy (anonymous):

ok so its the car + Tax

OpenStudy (anonymous):

I dont know why i really dont get your answer at all....8648+8648*.047=8648????

OpenStudy (anonymous):

Tax is basically a percentage of how much the cat costs SO the car costs 8646 and you pay 4.7% tax off of that So tax is 4,7 % of the 8648 so tax =8648*.047=? So you follow that?

OpenStudy (anonymous):

okay that equals $406.47?

OpenStudy (anonymous):

do you add that to the cost of the car and then start on the 5% part?

OpenStudy (anonymous):

yes ok but that is our tax SOOOOO Total cost = Cost of Car + Tax Total cost = 8648 + (8648*.047)

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

DO you see how i got my equation>

OpenStudy (anonymous):

NO i dont.... i dont see why you are adding the car cost twice? total cost= 8,648+406.47= 9,054.47?

OpenStudy (anonymous):

K cuz they wanna know how much they paid in total The paid for the car PLUSSSS they paid taxes

OpenStudy (anonymous):

So obviously he paid for his car and he also paid for taxes So his total cost are the cost of the car +taxes

OpenStudy (anonymous):

yeah car+tax is 8,648+406.47= 9,054.47

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

9,054.47*.05% ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i got 4.53 as an answer to that?

OpenStudy (anonymous):

$ 453 yea

OpenStudy (anonymous):

4 dollars. really.

OpenStudy (anonymous):

or is it $453.00??

OpenStudy (anonymous):

453.00

OpenStudy (anonymous):

k dont have time to help you with the next question

OpenStudy (anonymous):

repost it

OpenStudy (anonymous):

then why did it come out on my calculator as 4 dollars? okay and thank you